# The Analysis of Willmore Surfaces and its Generalizations in Higher Dimensions

### v1

@@ IV.4.2 Convergence of conformal structure
Let $\chi(\Sigma)$ denote the Euler characteristic of $\Sigma$, $K_{h_{k}}$ be the Gaussian curvature of $h_{k}$.
(Case I) $\chi(\Sigma)\leq-2\Rightarrow K_{h_{k}}=-1$.
By a theorem of Poincaré and its corollary (see [185, Theorem 1.6.2]), in this case, for any $s\in\mathbb{N}$, $s\geq 3$ and any metric $g\in W^{s,2}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$, there exists a unique metric $h\in W^{s,2}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$ such that $h$ is weakly conformal to $g$ and $h$ has constant Gaussian curvature 1-1. Moreover, the map taking $g$ to $h$ is continuous in $W^{s,2}$. It follows that $h_{k}$ is uniquely determined by $g_{k}$, and converges to $h_{\infty}$ smoothly since $g_{k}\rightarrow g_{\infty}$ in $C^{\infty}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$.
(Case II) $\chi(\Sigma)=0\Rightarrow K_{h_{k}}=0$.
Let $h_{k}=e^{2\lambda_{k}}g_{k}$ for some $\lambda_{k}\in C^{\infty}(\Sigma)$. We denote the Gaussian curvature and negative Laplace–Beltrami operator associated to $g_{k}$ by $K_{g_{k}}$ and $\Delta_{g_{k}}$ respectively, the condition that $h_{k}$ has constant Gaussian curvature 0 is then equivalent to the equation (see e.g. [168, Theorem 2.6])
$\Delta_{g_{k}}\lambda_{k}=K_{g_{k}}.$ (IV.57)
@@
Hence for each $k\in\mathbb{N}\cup\{\infty\}$, there exists a smooth solution $\lambda_{k}$ to (IV.57) and it is unique up to additive constants (see [9, Theorem 4.7]). By adding suitable constants, we can without loss of generality assume that
$\int\lambda_{k}\,d\textup{vol}_{g_{\infty}}=0\quad\text{for any }k\in\mathbb{N}\cup\{\infty\}.$ (IV.58)
Since $g_{k}\rightarrow g_{\infty}$ in $C^{\infty}$, we have $K_{g_{k}}\rightarrow K_{g_{\infty}}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$, and in particular, $\{K_{g_{k}}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $C^{0}(\Sigma)$. By (IV.57) and the convergence $g_{k}\rightarrow g_{\infty}$, there exists a constant $C_{1}>0$ independent of $k$ such that
$C_{1}\int_{\Sigma}|d\lambda_{k}|_{g_{\infty}}^{2}\,d\textup{vol}_{g_{\infty}}$ $\leq\int_{\Sigma}|d\lambda_{k}|_{g_{k}}^{2}\,d\textup{vol}_{g_{k}}$
$=-\int_{\Sigma}\lambda_{k}\,\Delta_{g_{k}}\lambda_{k}\,d\textup{vol}_{g_{k}}$
$=\int_{\Sigma}-\lambda_{k}K_{g_{k}}\,d\textup{vol}_{g_{k}}.$
Furthermore, by (IV.58), Cauchy–Schwarz inequality, and Poincaré’s inequality, there exist constants $C_{2},C_{3}>0$ independent of $k$ such that
$\int_{\Sigma}-\lambda_{k}K_{g_{k}}\,d\textup{vol}_{g_{k}}$ $\leq\left(\int_{\Sigma}\lambda_{k}^{2}\,d\textup{vol}_{g_{k}}\right)^{\frac{1}{2}}\left(\int_{\Sigma}K_{g_{k}}^{2}\,d\textup{vol}_{g_{k}}\right)^{\frac{1}{2}}$
$\leq C_{2}\left(\int_{\Sigma}\lambda_{k}^{2}\,d\textup{vol}_{g_{\infty}}\right)^{\frac{1}{2}}$
@@
Consequently, $\{\lambda_{k}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $W^{1,2}(\Sigma)$. Now we apply the interior estimates as in [58, Section 6.3.1] to the equation (IV.57) and obtain that $\{\lambda_{k}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $W^{s,2}(\Sigma)$ for any $s\in\mathbb{N}$. A standard compactness argument then implies $\lambda_{k}\rightarrow\lambda_{\infty}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$. Finally, by setting $h_{k}=e^{2\lambda_{k}}g_{k}$, we obtain the required convergent sequence of metrics of Gaussian curvature 0.
(Case III) $\chi(\Sigma)=2\Rightarrow K_{h_{k}}=1$.
@@
where $dz=dx+idy$ and $d\bar{z}=dx-idy$. Moreover, such $\mu_{k}$ is uniquely determined by $g_{k}$ and can be computed as (see e.g. [152, Section 1.2.1])
@@
In particular, $f_{k}$ is quasiconformal with complex dilatation $\mu_{k}$ (see Definition III.6). By considering the quasiconformal maps $f_{k}\circ f_{\infty}^{-1}$ and applying [90, Proposition 4.36], we have $f_{k}\rightarrow f_{\infty}$ locally uniformly on $\mathbb{C}$. Define $\sigma_{k}\in L^{\infty}(\mathbb{C},\mathbb{R}^{2\times 2})$ be the unique matrix-valued function on $\mathbb{C}$ such that $\sigma_{k}$ is positive definite a.e., $\text{det}(\sigma_{k})=1$, and
@@
Such $\sigma_{k}$ depends uniquely and smoothly on $\mu_{k}$, hence $\sigma_{k}\rightarrow\sigma_{\infty}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C},\mathbb{R}^{2\times 2})$. By [163, Theorem 5.1], $f_{k}$ satisfies the elliptic equation
@@
Hence standard elliptic estimates as in Case II imply that $f_{k}\rightarrow f_{\infty}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C})$. On the other hand, note $\psi\circ f_{k}\circ\psi^{-1}$ is quasiconformal with complex dilatation $\mu_{k}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}$ (see [90, Proposition 4.13]), and
@@
By the same argument as before, we have
@@
Hence $f_{k}\rightarrow f_{\infty}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$. Besides, by the proof of [4, Lemma 5.B.3], both $f_{k}$ and $\psi\circ f_{k}\circ\psi^{-1}$ have no critical points on $\mathbb{C}$, hence each $f_{k}$ is a diffeomorphism on $\Sigma$. Finally, for all $k\in\mathbb{N}\cup\{\infty\}$, let $h_{k}\coloneqq f_{k}^{*}h$, then $h_{k}$ is a Riemannian metric with constant Gaussian curvature $1$ on $\Sigma$, and $h_{k}\rightarrow h_{\infty}$ smoothly on $\Sigma$. Since $h$ is conformal to $g_{\text{std}}$ and $f_{k}^{*}g_{\text{std}}$ is conformal to $g_{k}$ on $\mathbb{C}$ (see for instance [90, Section 1.5.1]), we have $h_{k}=f_{k}^{*}h$ is conformal to $g_{k}$ on $\mathbb{C}\subset\Sigma$. The conformality holds on $\Sigma$ since $h_{k}$, $g_{k}$ are both Riemannian metrics on $\Sigma$, and $\mathbb{C}$ is dense in $\Sigma$. ∎
Then we consider the general case in which $\{g_{k}\}$ are metrics induced by weak immersions. The strategy of proof is an adaptation of [186, Proposition 3.2]. For a sequence of functions $f_{k}$ and a compact set $K\subset\mathbb{R}^{n}$, we write $f_{k}\rightarrow f$ in $C^{\infty}(K)$ if each $f_{k}$ is defined on a neighborhood of $K$ (possibly depending on $k$) and for any $l\in\mathbb{N}_{0}$, $\|f_{k}-f\|_{C^{l}(K)}\rightarrow 0$ as $k\rightarrow\infty$.
@@
Let $\Sigma$ be an oriented connected $2$-dimensional closed smooth manifold with a reference Riemannian metric $\tilde{g}$. Assume $\vec{\Phi}_{k}\rightarrow\vec{\Phi}_{\infty}$ in $W^{2,2}(\Sigma,\mathbb{R}^{m})$ as $k\rightarrow\infty$, and there exists a constant $\Lambda>0$ such that for any $k\in\mathbb{N}\cup\{\infty\}$, there holds
$\Lambda^{-1}|X|_{\tilde{g}}\leq|d(\vec{\Phi}_{k})_{p}(X)|_{\mathbb{R}^{m}}\leq\Lambda|X|_{\tilde{g}},\qquad\text{for a.e. }p\in\Sigma\text{ and all }\mkern 2.0muX\in T_{p}\Sigma.$ (IV.59)
Let $g_{k}\coloneqq\vec{\Phi}_{k}^{*}\,g_{\text{std}}$ denote the metric induced by $\vec{\Phi}_{k}$. Assume $\tilde{g}$ and $g_{\infty}$ are weakly conformal (i.e. $\Sigma$ is equipped with the complex structure induced by $\vec{\Phi}_{\infty}$ as in Corollary IV.5), then there exists a sequence $(h_{k})_{k\in\mathbb{N}\cup\{\infty\}}$ of metrics smooth in the complex structure induced by $\vec{\Phi}_{k}$, and a sequence $(\Psi_{k})_{k\in\mathbb{N}}$ of bi-Lipschitz homeomorphisms on $\Sigma$ such that
@@
$\varphi_{k,\alpha,\beta}\coloneqq\varphi_{k,\alpha}\circ\varphi_{k,\beta}^{-1}\xrightarrow[k\rightarrow\infty]{}\varphi_{\infty,\alpha}\circ\varphi_{\infty,\beta}^{-1}=:\varphi_{\infty,\alpha,\beta}\quad\text{in $W^{2,2}(K)$.}$ (IV.60)
@@
$\begin{cases}(i)\ \rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id&\text{ in }C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha}}),\mathbb{R}^{2}),\\[8.53581pt] (ii)\ \varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,\beta}=\rho_{k,\beta}&\text{ on }\varphi_{\infty,\beta}(\overline{V_{\alpha}}\cap\overline{V_{\beta}}).\end{cases}$ (IV.61)
@@
Once these maps $\rho_{k,\alpha}$ are defined, we consider the map $\Psi_{k}\coloneqq\varphi_{k,\alpha}^{-1}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha}$ on $\overline{V_{\alpha}}$. This definition is independent of $\alpha$ by (IV.61). By construction, it holds $\Psi_{k}\rightarrow id$ in $C^{0}(\Sigma,\Sigma)$. For $k$ large enough, $\Psi_{k}$ is an immersion, hence surjective (its image is both open and closed). As for global injectivity, assume there exist sequences $\{k_{i}\}_{i=1}^{\infty}\subset\mathbb{N}$, $\{p_{i,1}\}_{i=1}^{\infty}$, $\{p_{i,2}\}_{i=1}^{\infty}\subset\Sigma$ such that $\lim_{i\rightarrow\infty}k_{i}=\infty$, $\Psi_{k_{i}}(p_{i,1})=\Psi_{k_{i}}(p_{i,2})$. We can without loss of generality assume that $p_{i,1}\rightarrow p_{1}$ and $p_{i,2}\rightarrow p_{2}$ as $i\rightarrow\infty$, but we have $\Psi_{k}\rightarrow id$ in $C^{0}(\Sigma,\Sigma)$ as $k\rightarrow\infty$, hence $p_{1}=p_{2}\in V_{\alpha}$ for some $\alpha$. For $i$ large enough we get $p_{i,1}=p_{i,2}$ since $\Psi_{k}$ is injective on $V_{\alpha}$.
@@
We now proceed to the construction of $\{V_{\alpha}\}$ and $\{\rho_{k,\alpha}\}$. Let $\{U_{\alpha}^{\prime}\}_{\alpha=1}^{l}$ be an open cover of $\Sigma$ such that $U_{\alpha}^{\prime}$ has smooth boundary and $U_{\alpha}^{\prime}\Subset U_{\alpha}$ (for existence of such sets, see the proof of [52, Proposition 8.2.1]). We claim that there exist open sets $\{V_{\alpha,j}\}_{1\leq\alpha\leq j\leq l}$ (write $V_{\alpha,\alpha-1}=V_{\alpha,\alpha}$), $k_{0}\in\mathbb{N}$, and $C^{\infty}$ injective maps $\rho_{k,\alpha}\colon\varphi_{\infty,\alpha}(\overline{V_{\alpha,\alpha}})\rightarrow\varphi_{k,\alpha}(U_{\alpha})$ for $k\geq k_{0}$ such that for any $1\leq j\leq l$ and $1\leq\alpha,\beta\leq j$, it holds
$(i)\ V_{\alpha,j}\subset V_{\alpha,j-1}\subset U_{\alpha}^{\prime},$ (IV.62a)
$(ii)\ \Sigma=\Big(\bigcup\limits_{\gamma\leq j}V_{\gamma,j}\Big)\cup\Big(\bigcup\limits_{\gamma>j}U^{\prime}_{\gamma}\Big),$ (IV.62b)
$(iii)\ \rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id$ in $C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha,\alpha}}),\mathbb{R}^{2})$, (IV.62c)
$(iv)\ \varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,\beta}=\rho_{k,\beta}$ on $\varphi_{\infty,\beta}(\overline{V_{\alpha,j}}\cap\overline{V_{\beta,j}})$. (IV.62d)
@@
The proof is by induction on $j$. For $j=1$, we define $V_{1,1}=U_{1}^{\prime}$ and $\rho_{k,1}\colon\varphi_{\infty,1}(\overline{U_{1}^{\prime}})\rightarrow\mathbb{R}^{2}$ as the identity map. For $k$ large enough, the image of $\rho_{k,1}$ is contained in $\varphi_{k,1}(U_{1})$ since $\varphi_{k,1}\rightarrow\varphi_{\infty,1}$ in $C(\overline{U_{1}^{\prime}},\mathbb{R}^{2})$ and $U_{1}^{\prime}\Subset U_{1}$. Let $1\leq j_{0}\leq l$, suppose that we have constructed $\{V_{\alpha,j}\}_{1\leq\alpha\leq j\leq j_{0}}$ and $\{\rho_{k,\alpha}\}_{1\leq\alpha\leq j_{0}}$ such that (IV.62a)–(IV.62d) hold for any $1\leq j\leq j_{0}$ and $1\leq\alpha,\beta\leq j$. We now construct $\rho_{k,j_{0}+1}$ and $\{V_{\alpha,j_{0}+1}\}_{1\leq\alpha\leq j_{0}+1}$ using the induction hypothesis.
@@
By our induction hypothesis (IV.62d) for $j\leq j_{0}$, this map is well-defined on $\varphi_{\infty,j_{0}+1}\Big(\big(\bigcup\limits_{\alpha\leq j_{0}}\overline{V_{\alpha,j_{0}}}\big)\cap U_{j_{0}+1}\Big)$ and independent of $\alpha$. By the assumption $\Sigma=\big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\big)\cup\big(\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\big)$, we have
@@
The condition (IV.62a) then holds for $j=j_{0}+1$. By the definition of $\{V_{\alpha,j_{0}+1}\}_{\alpha\leq j_{0}+1}$ we have

### v3

@@ IV.4.2 Convergence of conformal structures
Let $\chi(\Sigma)$ denote the Euler characteristic of $\Sigma$ and $K_{h_{k}}$ be the Gaussian curvature of $h_{k}$.
(Case I) $\chi(\Sigma)\leq-2$ and $K_{h_{k}}=-1$.
By a theorem of Poincaré and its corollary (see [188, Theorem 1.6.2]), in this case, for any $s\in\mathbb{N}$, $s\geq 3$ and any metric $g\in W^{s,2}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$, there exists a unique metric $h\in W^{s,2}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$ such that $h$ is weakly conformal to $g$ and $h$ has constant Gaussian curvature 1-1. Moreover, the map taking $g$ to $h$ is continuous in $W^{s,2}$. It follows that $h_{k}$ is uniquely determined by $g_{k}$, and converges to $h_{\infty}$ smoothly since $g_{k}\rightarrow g_{\infty}$ in $C^{\infty}(\Sigma,T^{*}\Sigma\otimes T^{*}\Sigma)$.
(Case II) $\chi(\Sigma)=0$ and $K_{h_{k}}=0$.
Let $h_{k}=e^{2\lambda_{k}}g_{k}$ for some $\lambda_{k}\in C^{\infty}(\Sigma)$. We denote the Gaussian curvature and negative Laplace–Beltrami operator associated to $g_{k}$ by $K_{g_{k}}$ and $\Delta_{g_{k}}$ respectively, the condition that $h_{k}$ has constant Gaussian curvature 0 is then equivalent to the equation (see e.g. [171, Theorem 2.6])
$\Delta_{g_{k}}\lambda_{k}=K_{g_{k}}.$ (IV.58)
@@
Hence for each $k\in\mathbb{N}\cup\{\infty\}$, there exists a smooth solution $\lambda_{k}$ to (IV.58) and it is unique up to additive constants (see [9, Theorem 4.7]). By adding suitable constants, we can without loss of generality assume that
$\forall k\in\mathbb{N}\cup\{\infty\},\qquad\int_{\Sigma}\lambda_{k}\,d\textup{vol}_{g_{\infty}}=0.$ (IV.59)
Since $g_{k}\rightarrow g_{\infty}$ in $C^{\infty}$, we have $K_{g_{k}}\rightarrow K_{g_{\infty}}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$, and in particular, $\{K_{g_{k}}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $C^{0}(\Sigma)$. By (IV.58) and the convergence $g_{k}\rightarrow g_{\infty}$, there exists a constant $C_{1}>0$ independent of $k$ such that
$C_{1}\int_{\Sigma}|d\lambda_{k}|_{g_{\infty}}^{2}\,d\textup{vol}_{g_{\infty}}$ $\leq\int_{\Sigma}|d\lambda_{k}|_{g_{k}}^{2}\,d\textup{vol}_{g_{k}}=-\int_{\Sigma}\lambda_{k}\,\Delta_{g_{k}}\lambda_{k}\,d\textup{vol}_{g_{k}}=\int_{\Sigma}-\lambda_{k}K_{g_{k}}\,d\textup{vol}_{g_{k}}.$
Furthermore, by (IV.59), Cauchy–Schwarz inequality, and Poincaré inequality, there exist constants $C_{2},C_{3}>0$ independent of $k$ such that
$\int_{\Sigma}-\lambda_{k}K_{g_{k}}\,d\textup{vol}_{g_{k}}$ $\leq\|K_{g_{k}}\|_{L^{\infty}(\Sigma)}\,\textup{vol}_{g_{k}}(\Sigma)^{\frac{1}{2}}\left(\int_{\Sigma}\lambda_{k}^{2}\,d\textup{vol}_{g_{k}}\right)^{\frac{1}{2}}$
$\leq C_{2}\,\left(\sup_{l\in\mathbb{N}}\|K_{g_{l}}\|_{L^{\infty}(\Sigma)}\right)\,\left(\sup_{l\in\mathbb{N}}\textup{vol}_{g_{l}}(\Sigma)^{\frac{1}{2}}\right)\left(\int_{\Sigma}\lambda_{k}^{2}\,d\textup{vol}_{g_{\infty}}\right)^{\frac{1}{2}}$
@@
Consequently, $\{\lambda_{k}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $W^{1,2}(\Sigma)$. Now we apply the interior estimates as in [58, Section 6.3.1] to the equation (IV.58) and obtain that $\{\lambda_{k}\}_{k\in\mathbb{N}\cup\{\infty\}}$ is bounded in $W^{s,2}(\Sigma)$ for any $s\in\mathbb{N}$. A standard compactness argument then implies $\lambda_{k}\rightarrow\lambda_{\infty}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$. Finally, by setting $h_{k}=e^{2\lambda_{k}}g_{k}$, we obtain the required convergent sequence of metrics of Gaussian curvature 0.
(Case III) $\chi(\Sigma)=2$ and $K_{h_{k}}=1$.
@@
where $dz=dx+idy$ and $d\bar{z}=dx-idy$. Moreover, such $\mu_{k}$ is uniquely determined by $g_{k}$ and can be computed as (see e.g. [155, Section 1.2.1])
@@
In particular, $f_{k}$ is quasiconformal with complex dilatation $\mu_{k}$ (see Definition III.6). By considering the quasiconformal maps $f_{k}\circ f_{\infty}^{-1}$ and applying [91, Proposition 4.36], we have $f_{k}\rightarrow f_{\infty}$ locally uniformly on $\mathbb{C}$. We denote $\sigma_{k}\in L^{\infty}(\mathbb{C},\mathbb{R}^{2\times 2})$ the unique matrix-valued function on $\mathbb{C}$ such that $\sigma_{k}$ is positive definite a.e., $\text{det}(\sigma_{k})=1$, and such that
@@
Such $\sigma_{k}$ depends uniquely and smoothly on $\mu_{k}$, hence we have $\sigma_{k}\rightarrow\sigma_{\infty}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C},\mathbb{R}^{2\times 2})$. By [166, Theorem 5.1], $f_{k}$ satisfies the elliptic equation
@@
Hence standard elliptic estimates as in Case II imply that $f_{k}\rightarrow f_{\infty}$ in $C^{\infty}_{\textup{loc}}(\mathbb{C})$. On the other hand, the map $\psi\circ f_{k}\circ\psi^{-1}$ is quasiconformal with complex dilatation $\mu_{k}\circ\psi^{-1}(\omega)\,\omega^{2}{\bar{\omega}}^{-2}$ (see [91, Proposition 4.13]). Moreover, it holds
@@
By the same argument as before, we obtain
@@
Hence it holds $f_{k}\rightarrow f_{\infty}$ in $C^{\infty}(\Sigma)$ as $k\rightarrow\infty$. Besides, by the proof of [4, Lemma 5.B.3], both $f_{k}$ and $\psi\circ f_{k}\circ\psi^{-1}$ have no critical points on $\mathbb{C}$, hence each $f_{k}$ is a diffeomorphism on $\Sigma$. Finally, for all $k\in\mathbb{N}\cup\{\infty\}$, let $h_{k}\coloneqq f_{k}^{*}h$, then $h_{k}$ is a Riemannian metric with constant Gaussian curvature $1$ on $\Sigma$, and $h_{k}\rightarrow h_{\infty}$ smoothly on $\Sigma$. Since $h$ is conformal to $g_{\text{std}}$ and $f_{k}^{*}g_{\text{std}}$ is conformal to $g_{k}$ on $\mathbb{C}$ (see for instance [91, Section 1.5.1]), we have $h_{k}=f_{k}^{*}h$ is conformal to $g_{k}$ on $\mathbb{C}\subset\Sigma$. The conformality holds on $\Sigma$ since $h_{k}$, $g_{k}$ are both Riemannian metrics on $\Sigma$, and $\mathbb{C}$ is dense in $\Sigma$. ∎
Then we consider the general case in which $\{g_{k}\}$ are metrics induced by weak immersions. The strategy of proof is an adaptation of [189, Proposition 3.2]. For a sequence of functions $f_{k}$ and a compact set $K\subset\mathbb{R}^{n}$, we write $f_{k}\rightarrow f$ in $C^{\infty}(K)$ if each $f_{k}$ is defined on a neighborhood of $K$ (possibly depending on $k$) and for any $l\in\mathbb{N}_{0}$, $\|f_{k}-f\|_{C^{l}(K)}\rightarrow 0$ as $k\rightarrow\infty$.
@@
Let $\Sigma$ be an oriented connected $2$-dimensional closed smooth manifold with a reference Riemannian metric $\tilde{g}$. Assume that $\vec{\Phi}_{k}\rightarrow\vec{\Phi}_{\infty}$ in $W^{2,2}(\Sigma,\mathbb{R}^{m})$ as $k\rightarrow\infty$, and that there exists a constant $\Lambda>0$ such that for any $k\in\mathbb{N}\cup\{\infty\}$, there holds
$\Lambda^{-1}|X|_{\tilde{g}}\leq|d(\vec{\Phi}_{k})_{p}(X)|_{\mathbb{R}^{m}}\leq\Lambda|X|_{\tilde{g}},\qquad\text{for a.e. }p\in\Sigma\text{ and all }\mkern 2.0muX\in T_{p}\Sigma.$ (IV.60)
Let $g_{k}\coloneqq\vec{\Phi}_{k}^{*}\,g_{\text{std}}$ be the metric induced by $\vec{\Phi}_{k}$. Assume that $\tilde{g}$ and $g_{\infty}$ are weakly conformal (i.e. $\Sigma$ is equipped with the complex structure induced by $\vec{\Phi}_{\infty}$ as in Corollary IV.5), then there exists a sequence $(h_{k})_{k\in\mathbb{N}\cup\{\infty\}}$ of metrics smooth in the complex structure induced by $\vec{\Phi}_{k}$, and a sequence $(\Psi_{k})_{k\in\mathbb{N}}$ of bi-Lipschitz homeomorphisms on $\Sigma$ such that
@@
$\varphi_{k,\alpha,\beta}\coloneqq\varphi_{k,\alpha}\circ\varphi_{k,\beta}^{-1}\xrightarrow[k\rightarrow\infty]{}\varphi_{\infty,\alpha}\circ\varphi_{\infty,\beta}^{-1}=:\varphi_{\infty,\alpha,\beta}\quad\text{in $W^{2,2}(K)$.}$ (IV.61)
@@
$\begin{cases}(i)\ \rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id&\text{ in }C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha}}),\mathbb{R}^{2}),\\[8.53581pt] (ii)\ \varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,\beta}=\rho_{k,\beta}&\text{ on }\varphi_{\infty,\beta}(\overline{V_{\alpha}}\cap\overline{V_{\beta}}).\end{cases}$ (IV.62)
@@
Once these maps $\rho_{k,\alpha}$ are defined, we consider the map $\Psi_{k}\coloneqq\varphi_{k,\alpha}^{-1}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha}$ on $\overline{V_{\alpha}}$. This definition is independent of $\alpha$ by (IV.62). By construction, it holds $\Psi_{k}\rightarrow id$ in $C^{0}(\Sigma,\Sigma)$. For $k$ large enough, $\Psi_{k}$ is an immersion, hence surjective (its image is both open and closed). As for global injectivity, assume that there exist sequences $\{k_{i}\}_{i=1}^{\infty}\subset\mathbb{N}$, $\{p_{i,1}\}_{i=1}^{\infty}$, $\{p_{i,2}\}_{i=1}^{\infty}\subset\Sigma$ such that $\lim_{i\rightarrow\infty}k_{i}=\infty$ and $\Psi_{k_{i}}(p_{i,1})=\Psi_{k_{i}}(p_{i,2})$. We can without loss of generality assume that $p_{i,1}\rightarrow p_{1}$ and $p_{i,2}\rightarrow p_{2}$ as $i\rightarrow\infty$. Since we have $\Psi_{k}\rightarrow id$ in $C^{0}(\Sigma,\Sigma)$ as $k\rightarrow\infty$, we obtain $p_{1}=p_{2}\in V_{\alpha}$ for some $\alpha$. For $i$ large enough we get $p_{i,1}=p_{i,2}$ since $\Psi_{k}$ is injective on $V_{\alpha}$.
@@
We now proceed to the construction of $\{V_{\alpha}\}$ and $\{\rho_{k,\alpha}\}$. Let $\{U_{\alpha}^{\prime}\}_{\alpha=1}^{l}$ be an open cover of $\Sigma$ such that $U_{\alpha}^{\prime}$ has smooth boundary and $U_{\alpha}^{\prime}\Subset U_{\alpha}$ (for existence of such sets, see the proof of [52, Proposition 8.2.1]). We claim that there exist open sets $\{V_{\alpha,j}\}_{1\leq\alpha\leq j\leq l}$ (write $V_{\alpha,\alpha-1}=V_{\alpha,\alpha}$), an integer $k_{0}\in\mathbb{N}$, and $C^{\infty}$ injective maps $\rho_{k,\alpha}\colon\varphi_{\infty,\alpha}(\overline{V_{\alpha,\alpha}})\rightarrow\varphi_{k,\alpha}(U_{\alpha})$ for $k\geq k_{0}$ such that for any $1\leq j\leq l$ and $1\leq\alpha,\beta\leq j$, it holds
$(i)\ V_{\alpha,j}\subset V_{\alpha,j-1}\subset U_{\alpha}^{\prime},$ (IV.63a)
$(ii)\ \Sigma=\Big(\bigcup\limits_{\gamma\leq j}V_{\gamma,j}\Big)\cup\Big(\bigcup\limits_{\gamma>j}U^{\prime}_{\gamma}\Big),$ (IV.63b)
$(iii)\ \rho_{k,\alpha}\xrightarrow[k\rightarrow\infty]{}id$ in $C^{\infty}(\varphi_{\infty,\alpha}(\overline{V_{\alpha,\alpha}}),\mathbb{R}^{2})$, (IV.63c)
$(iv)\ \varphi_{k,\beta,\alpha}\circ\rho_{k,\alpha}\circ\varphi_{\infty,\alpha,\beta}=\rho_{k,\beta}$ on $\varphi_{\infty,\beta}(\overline{V_{\alpha,j}}\cap\overline{V_{\beta,j}})$. (IV.63d)
@@
We prove this by induction on $j$. For $j=1$, we define $V_{1,1}=U_{1}^{\prime}$ and $\rho_{k,1}\colon\varphi_{\infty,1}(\overline{U_{1}^{\prime}})\rightarrow\mathbb{R}^{2}$ as the identity map. For $k$ large enough, the image of $\rho_{k,1}$ is contained in $\varphi_{k,1}(U_{1})$ since $\varphi_{k,1}\rightarrow\varphi_{\infty,1}$ in $C(\overline{U_{1}^{\prime}},\mathbb{R}^{2})$ and $U_{1}^{\prime}\Subset U_{1}$. Let $1\leq j_{0}\leq l$, suppose that we have constructed $\{V_{\alpha,j}\}_{1\leq\alpha\leq j\leq j_{0}}$ and $\{\rho_{k,\alpha}\}_{1\leq\alpha\leq j_{0}}$ such that (IV.63a)–(IV.63d) hold for any $1\leq j\leq j_{0}$ and $1\leq\alpha,\beta\leq j$. We now construct $\rho_{k,j_{0}+1}$ and $\{V_{\alpha,j_{0}+1}\}_{1\leq\alpha\leq j_{0}+1}$ using the induction hypothesis.
@@
By our induction hypothesis (IV.63d) for $j\leq j_{0}$, this map is well-defined on $\varphi_{\infty,j_{0}+1}\Big(\big(\bigcup\limits_{\alpha\leq j_{0}}\overline{V_{\alpha,j_{0}}}\big)\cap U_{j_{0}+1}\Big)$ and independent of $\alpha$. By the assumption $\Sigma=\big(\bigcup\limits_{\alpha\leq j_{0}}V_{\alpha,j_{0}}\big)\cup\big(\bigcup\limits_{\alpha>j_{0}}U^{\prime}_{\alpha}\big)$, we have
@@
The condition (IV.63a) then holds for $j=j_{0}+1$. By the definition of $\{V_{\alpha,j_{0}+1}\}_{\alpha\leq j_{0}+1}$ we have